OpenStudy (anonymous):

the electric force between two charged objects is .24 newtons. If the distance between them is increased by a factor of 4, what will be the new force

OpenStudy (anonymous):

F1 = k q Q/ r^2 = 0.24 N F2 = k q Q/ (4r)^2 F2/F1 = (r)^2/ (4r)^2 and the rest is yours